# Indoor Location Fingerprinting Privacy: A Comprehensive Survey

### v1

@@ 4.1.2. Users
The user itself can also act as a malicious party. Malicious users’ actions in indoor localization are twofold: 1) Malicious users can access unauthorized information on the server through security misconfiguration and hacking or capture other users’ data to track other users or sell the information for profit. The localization server and localization process should protect this data from being exposed; 2) Collision of malicious users is another threat. Peers subscribing to the same LBS can either collude to launch attacks, or one adversary can create fake peers to obtain the information they seek. For example by creating three fake anchor locations and using their corresponding distances to the target in an iterative trilateration based on a localization algorithm to obtain an inferred location (Liu et al., 2018).
4.1.3. Trusted Third Party (TTP)
@@ 4.1.4. LSP
4.1.5. Content Provider
Some contents, including maps and POIs, are essential for LBS functionality. While some LBS providers employ a third party to deliver this service, others use their own content. For indoor localization, the content may include, the building map and semantic information of each section or room. Privacy of this content is also important as it has information about different parts of the buildings and the service provider would not like it to be publicly available, mostly for sensitive data. Consider a hospital where the service provider is reluctant for users to see the sensitive parts of the building, such as the electrical facilities, clean rooms, etc.
@@ 4.1.6. Data Publisher
The attacker can link the data of the published dataset with other public datasets to identify or re-identify users or other unauthorized activities. Similar to the Content Provider itself cannot introduce a privacy threat to the IPS, as the Data Publisher’s communication with the server is unilateral.
4.2. Data Structure-Driven Privacy Leakage
In this section, inherent privacy vulnerabilities in the data structure of the indoor location fingerprinting systems are introduced. The location data in LBSs varies in indoor environments due to the different methods used for indoor localization. Different from outdoor positioning systems, where the main data is the coordinates (e.g. longitude and latitude), the data for indoor environments employing the fingerprinting technique is various signal characteristics vectors. As shown in Fig. 5, the data in indoor location fingerprinting transmitted through the network as the tuple of $<ID,Time,location>$ (Wernke et al., 2014). The $ID$ and $Time$ data can be defined the same for indoor and outdoor localization, but the $location$ data has a different type. The terms are explained below:
Figure 5. Data structure in indoor location fingerprinting systems.
4.2.1. ID
Identity refers to a user’s name, email address, or any other characteristic that distinguishes one individual from another. Identifications in LBSs can either be consistent or inconsistent (Zheng and Zhou, 2011). Certain LBSs require constant user IDs, while others do not necessitate a user’s identity or a consistent one. For instance, users can employ LBSs under a fake name or in secret. The most common constant identity that LBSs require is an email address or a phone number. From another perspective, the ID can be hardcoded such as MAC-Address (or possibly IMEI). It can also be a costume ID given by software; for example, the session ID for data transmission or a self-defined ID is characterized by the application developer.
@@ 4.2. Data Structure-Driven Privacy Leakage
  • Single identity leaks or personal identification assaults involve pinpointing a user’s identity through their indoor location, such as within a faculty building’s lab, or indirectly by narrowing down identification possibilities—such as inferring gender, educational level, or access to restricted areas. This approach allows for further investigation to potentially uncover someone’s identity (Krumm, 2007). For instance, a solitary presence in a room, like a teacher’s office, offers a higher chance of exact identification, while a crowded space, like a student-filled lab, lowers the probability of specific identification but doesn’t eliminate it.
@@
  • Multiple identity information leaks, meeting disclosure leaks, or aggregated presence assaults try to determine the relationship between two or more people, or between people and an aggregated property, such as whether people meet at a particular time or the approximate number of persons visit a lab (Liu et al., 2018). As a case in point, when two or more people go to the same lab at specific time intervals, they are highly probably co-workers or friends, or they have a meeting at that time.
4.2.2. Location
@@
The primary tool for determining the location of an object is spatial information. Coordinates, such as longitude and latitude, are employed for outdoor locations, while signal characteristics are used for indoor localization. Other information, like a store’s name, may also be associated with a location. In indoor fingerprinting localization, hence, the location data involve the characteristics of the existing signal in the indoor area, such as RSS or CSI when RF technologies are utilized for localization, as mentioned in Section 2. This location data is received by the user’s device, sent to the server for the calculation of the user’s location, and the estimated location is then sent back to the user. Location-driven privacy leakage refers to locating crucial areas, such as a person’s exact workplace in an indoor environment (Gu et al., 2016).
4.2.3. Time
Some LBSs integrate identification and location information, along with a timestamp for each location. Regarding time, data transmission can be categorized as either real-time or non-real-time. Certain LBSs forego real-time services, opting to collect data for later publishes. Nevertheless, the primary objective for most LBSs is to provide real-time services, each posing its own set of privacy issues. Real-time privacy protection is more challenging to achieve than non-real-time protection due to the need for significantly more scalable solutions. Moreover, given the unpredictable and dynamic motions of users, achieving global optimization with real-time information is exceedingly challenging (Zheng and Zhou, 2011). Two examples of LBSs that utilize real-time location data are navigation and augmented reality games.
@@ 4.2.3. Time
Based on the combination of time and location data, the various types of spatial information can be broadly categorized into two groups: Single and Trajectory. Single location information refers to static spatial data that provides details about a specific point in the area. This category applies to situations where the focus is on a singular, fixed position, such as the layout of a room or the coordinates of a stationary object. In contrast, trajectory information involves the dynamic aspect of location data, capturing the movement and path of an object over time, such as tracking the trajectory of a person or an asset in indoor areas for navigation, asset management, or security purposes.
@@
4.3. Inference-Driven Privacy Leakage
@@
In indoor location fingerprinting, there are also some privacy leakages arising from the information inferred about the users (Hemkumar, 2024). Two critical aspects can be considered as the main dimensions that can be inferred and violate privacy in indoor location fingerprinting systems: identity inference and profiling completeness. Identity inference raises concerns about unauthorized tracking and monitoring, unveiling sensitive details through IPS. Simultaneously, profiling completeness, while enhancing user profiles, poses risks by holding sensitive information. These two privacy dimensions are explained below in the following subsections (Liu et al., 2018).
4.3.1. Identity inference
@@
Identity inference from indoor location data introduces multifaceted privacy concerns. Location tracking through IPS not only unveils sensitive details about a user’s activities, habits, and interests but also poses the risk of unauthorized tracking and monitoring without user consent (Krumm, 2007). The identification capabilities of these systems, based on unique location patterns, raise ethical concerns related to targeted marketing, personal profiling, and potential discriminatory practices. Furthermore, the prospect of re-identification, where seemingly anonymous users can be linked to their identity by combining location data with external information, underscores the need for stringent privacy safeguards. Additionally, location data can lead to potential inference that extends beyond exact location disclosure, as frequent visits to specific places can reveal significant information, such as health conditions. Collectively, these aspects highlight the intricate challenges associated with preserving identity privacy in the context of location data and emphasize the importance of ethical considerations and robust protective measures.
4.3.2. Profiling completeness
@@
Profiling completeness involves the frequent identification of Points of Interest (POIs) within a building based on user locations (Pombinho et al., 2011). Utilizing this information becomes crucial for gaining a deeper understanding of user behavior in real-world scenarios and predicting future actions. This data, commonly employed for marketing and surveillance purposes, possesses the capability to significantly enhance the comprehensiveness of profiles compared to basic data alone. The functionality of location data to seamlessly ”connect the dots” allows for the automated generation of profiles for individuals or organizations. IPS, pivotal for tracking a user’s position within a building, hold sensitive information. Unauthorized access to this information could lead to tracking and monitoring without user awareness or consent, echoing the challenges discussed in identity inference.
5. Attacks on Indoor Location Fingerprinting
@@
An adversary attacks IPS aiming to collect the location information of users and use it for their benefit. This section discusses the attacks on indoor location fingerprinting systems, investigating the different aspects including adversary and attack models to obtain meaningful data about the user’s identity, location, and time.
5.1. Adversary Models
@@
In general, user privacy means that no passive adversary (including a curious or malicious server) can determine the location of an honest user after intercepting all protocol messages. Server privacy, however, protects the server’s data from being compromised by malicious users using location queries (Yang and Järvinen, 2018). To discuss threat models on privacy preservation in indoor localization, we need to explain privacy assumptions on how entities can behave in IPS. Four assumptions are taken into consideration which are explained in the following subsection.
5.1.1. Fully trusted
@@
5.1.2. Semi-trusted (honest-but-curious)
@@ 5.1.3. Fully malicious
5.1.4. Uunilateral-malicious
Introduced in (Yang and Järvinen, 2018), the unilateral-malicious setting is weaker than the fully malicious setting but stronger than the conventional Semi-trusted setting. We formulate the malicious behaviors specific to user sessions in the unilateral-malicious setting, such as manipulating Wi-Fi fingerprints and disclosing locations. We expect the server to act fairly honestly. In other words, while the server may be interested in learning the location of a user, it should run the protocol instance honestly to offer a high-quality service. Since a server’s malicious actions, such as dishonest executions, would be easily detected in practice and punished due to providing inaccurate location services, we can lower the security requirements for the server. In other words, if the service is poor, then users would likely just stop using the service and, consequently, make such an attack impossible. However, the server cannot easily identify a user’s malicious behaviors.
Understanding the difference between various settings lies in the level of trust and cooperation expected from the participants, as well as the defences and security measures needed to ensure privacy and integrity in the face of potential adversarial behaviors and threats.
Figure 6. Adversary and attack models overview in in indoor location fingerprinting systems.
5.2. Attack Models
@@
5.2.1. Location Privacy and Data Privacy Attacks
Location privacy revolves around the potential threats to an individual’s privacy resulting from the collection, use, and disclosure of their location information. Targeted advertising, monitoring daily habits, surveillance, stalking and harassment, discriminatory uses, and unauthorized access are examples of location privacy concerns. Indoor locations due to more precise positioning are more susceptible to privacy issues. Overall, location privacy is a subcategory of data privacy that is disparate in importance. By this means, from the perspective of the user, different locations have different privacy requirements. For instance, the majority of people worry significantly more about keeping their houses and workplaces private than they do about revealing where they visited in a shopping center. Despite differing levels of data sensitivity, the dangers associated with location privacy remain a substantial concern (Liu et al., 2018).
The major attack model on IPS is introduced in (Li et al., 2014), which is followed by most papers. As shown in Fig. 7, this attack model categorizes the threats into the Location Privacy Attacks and Data Privacy Attacks. In both of them, location and data can be attacked in two ways as follows:
@@ 5.2.1. Location Privacy and Data Privacy Attacks
Figure 7. Location privacy and data privacy attacks in in indoor location fingerprinting systems.
@@
Note that another categorization with the same principle is applied by using terms user-side privacy attack and server-side privacy attack (Yang and Järvinen, 2018). The localization server and the APs are regarded as adversaries for preserving the user’s location privacy. The user, on the other hand, is viewed as an adversary to protect the server’s database. The user must not reveal her location to untrusted parties, and the service provider should also keep its WiFi fingerprint database from unauthorized leaking. For example, a malicious user may download the database and sell it for profit.
5.2.2. Protocol-based attack
@@
In this attack, the malicious actor uses cryptographic methods to extract data from encrypted data or discovers misconfiguration in data transmission protocols between user and server. Specifically in the chosen fingerprint attack (Yang and Järvinen, 2018) or known location attack (Yang and Järvinen, 2018), the user chooses special fingerprints, such as all-zeros or single-one fingerprints, to compromise the whole server’s database.
5.2.3. Inference attack
An inference attack is a type of attack in which an adversary deduces sensitive information about individuals or the system by analyzing seemingly innocuous or non-sensitive data. Inference attacks often exploit the relationships and patterns present in the available data to draw conclusions that may not be explicitly revealed. In the case of location-based services, inference attacks may attempt to deduce a person’s activities, habits, or even identity by analyzing patterns in the location data collected over time.
5.2.4. Semantic attack
@@
In a semantic attack, the objective of the attacker is to obtain specific semantic details about an individual. Location data inherently contains semantic information, and the disclosure of such information can unveil personal details and behavioral tendencies of users. Within indoor location datasets, an adversary can deduce the purpose of a user’s visits to particular locations by monitoring their movements or extended stays (Fathalizadeh et al., 2022). For example, in a hospital, the adversary may discern whether the user is visiting a specific doctor, and in a shopping center, they can identify specific shops where the user lingers. Additionally, this information might reveal whether the user is searching for or purchasing particular products in a shopping center, consulting a specific specialist in a hospital, or meeting a particular professor at a university.
5.2.5. Probability-based attack

### v2

@@ 3.1.2 Users
The user itself can also act as a malicious party. Malicious users’ actions in indoor localization are twofold: 1) Malicious users can access unauthorized information on the server through security misconfiguration and hacking or capture other users’ data to track other users or sell the information for profit. The localization server and localization process should protect this data from being exposed; 2) Collision of malicious users is another threat. Peers subscribing to the same LBS can either collude to launch attacks, or one adversary can create fake peers to obtain the information they seek. For example, malicious users might create three fake anchor locations and use their corresponding distances to the target in an iterative trilateration process, based on a localization algorithm, to infer a target’s location. This technique can mislead the localization system into producing inaccurate results or revealing sensitive information by exploiting the false data provided by these fake anchors.
3.1.3 Trusted Third Party (TTP)
@@ 3.1.4 LSP
3.1.5 Content Provider
Some contents, including maps and POIs, are essential for LBS functionality. While some LBS providers employ a third party to deliver this service, others use their own content. For indoor localization, the content may include the building map and semantic information of each section or room. Privacy of this content is also important as it has information about different parts of the buildings and the service provider would not like it to be publicly available, mostly for sensitive data. Consider a hospital where the service provider is reluctant for users to see the sensitive parts of the building, such as the electrical facilities, clean rooms, etc.
@@ 3.1.6 Data Publisher
The attacker can link the data of the published dataset with other public datasets to identify or re-identify users or other unauthorized activities. Similar to the Content Provider, Data Publishing itself cannot introduce a privacy threat to the IPS, as the Data Publisher’s communication with the server is unilateral.
3.2 Data Structure-Driven Privacy Leakage
In this section, inherent privacy vulnerabilities in the data structure of the ILF systems are introduced. The location data in LBSs varies in indoor environments due to the different methods used for indoor localization. Different from outdoor positioning systems, where the main data is the coordinates (e.g. longitude and latitude), the data for indoor environments employing the fingerprinting technique is various signal characteristics vectors. As shown in Fig. 5, the data in ILF is transmitted through the network as the tuple <ID,Time,location><ID,Time,location> [36]. The $ID$ and $Time$ data can be defined the same for indoor and outdoor localization, but the locationlocation data has a different type. The terms are explained below:
Figure 5: Data structure in ILF systems.
3.2.1 ID
Identity refers to a user’s name, email address, or any other characteristic that distinguishes one individual from another. In the context of LBSs, identities can be categorized as either consistent or inconsistent [37]. Consistent identities are those that remain the same across different sessions or interactions with the service, such as a fixed email address, phone number, or a hardcoded identifier like a MAC address or IMEI. These identifiers allow LBSs to consistently recognize and track the same user over time. Inconsistent identities, on the other hand, can vary between sessions or interactions. For instance, users might employ LBSs under a pseudonym or a temporary, self-defined ID provided by the application, such as a session ID or a custom identifier created for a specific purpose. This variability in identity helps protect user privacy by making it more challenging to link activities or track individuals across different sessions.
@@ 3.2 Data Structure-Driven Privacy Leakage
  • Single identity leaks or personal identification assaults involve pinpointing a user’s identity through their indoor location, such as within a faculty building’s lab, or indirectly by narrowing down identification possibilities—such as inferring gender, educational level, or access to restricted areas. This approach allows for further investigation to potentially uncover someone’s identity [38]. For instance, a solitary presence in a room, like a teacher’s office, offers a higher chance of exact identification, while a crowded space, like a student-filled lab, lowers the probability of specific identification but doesn’t eliminate it.
@@
  • Multiple identity information leaks, meeting disclosure leaks, or aggregated presence assaults try to determine the relationship between two or more people, or between people and an aggregated property, such as whether people meet at a particular time or the approximate number of persons visiting a lab [39]. As a case in point, when two or more people go to the same lab at specific time intervals, they are highly probably co-workers or friends, or they have a meeting at that time.
3.2.2 Location
@@
The primary tool for determining the location of an object is spatial information. Coordinates, such as longitude and latitude, are employed for outdoor locations, while signal characteristics are used for indoor localization. Other information, like a store’s name, may also be associated with a location. In indoor fingerprinting localization, hence, the location data involve the characteristics of the existing signal in the indoor area, such as RSS or CSI when RF technologies are utilized for localization, as mentioned in Section 2. This location data is received by the user’s device, sent to the server for the calculation of the user’s location, and the estimated location is then sent back to the user. Location-driven privacy leakage refers to locating crucial areas, such as a person’s exact workplace in an indoor environment [40].
3.2.3 Time
Some LBSs integrate identification and location information, along with a timestamp for each location. Data transmission can be categorized as either real-time or non-real-time. Certain LBSs opt for non-real-time data collection, saving information for later use, while others prioritize real-time services, each presenting unique privacy challenges. Real-time privacy protection is more complex due to the need for scalable solutions that can handle dynamic and unpredictable user movements. Achieving global optimization in this context refers to the challenge of effectively managing and analyzing real-time data across various locations and users to ensure accurate and efficient service delivery while maintaining privacy. This is particularly difficult because it requires balancing the need for precise, up-to-date information with robust privacy safeguards. Two examples of LBSs that utilize real-time location data are navigation systems and augmented reality games.
@@ 3.2.3 Time
Based on the combination of time and location data, spatial information can be broadly categorized into two groups: Single and Trajectory. Single location information refers to static spatial data about a specific point, such as the layout of a room or the coordinates of a stationary object. In contrast, trajectory information captures the movement and path of an object over time, used for tracking people or assets for navigation, asset management, or security purposes.
@@
3.3 Inference-Driven Privacy Leakage
@@
In ILF systems, there are also some privacy leakages arising from the information inferred about the users [41]. Two critical aspects can be considered as the main dimensions that can be inferred and violate privacy in indoor fingerprinting localization: identity inference and profiling completeness. Identity inference raises concerns about unauthorized tracking and monitoring, unveiling sensitive details through IPS. Simultaneously, profiling completeness, while enhancing user profiles, poses risks by holding sensitive information. These two privacy dimensions are explained below in the following subsections [39].
3.3.1 Identity inference
@@
Identity inference from indoor location data introduces multifaceted privacy concerns. Location tracking through IPS not only unveils sensitive details about a user’s activities, habits, and interests but also poses the risk of unauthorized tracking and monitoring without user consent [38]. The identification capabilities of these systems, based on unique location patterns, raise ethical concerns related to targeted marketing, personal profiling, and potential discriminatory practices. Furthermore, the prospect of re-identification, where seemingly anonymous users can be linked to their identity by combining location data with external information, underscores the need for stringent privacy safeguards. Additionally, location data can lead to potential inference that extends beyond exact location disclosure, as frequent visits to specific places can reveal significant information, such as health conditions. Collectively, these aspects highlight the intricate challenges associated with preserving identity privacy in the context of location data and emphasize the importance of ethical considerations and robust protective measures.
3.3.2 Profiling completeness
@@
Profiling completeness involves the frequent identification of Points of Interest (POIs) within a building based on user locations [42]. Utilizing this information becomes crucial for gaining a deeper understanding of user behavior in real-world scenarios and predicting future actions. This data, commonly employed for marketing and surveillance purposes, possesses the capability to significantly enhance the comprehensiveness of profiles compared to basic data alone. The functionality of location data to seamlessly ”connect the dots” allows for the automated generation of profiles for individuals or organizations. IPS, pivotal for tracking a user’s position within a building, holds sensitive information. Unauthorized access to this information could lead to tracking and monitoring without user awareness or consent, echoing the challenges discussed in identity inference.
4 Attacks on Indoor Location Fingerprinting
@@
An adversary attacks IPS aiming to collect the location information of users and use it for their benefit. This section discusses the attacks on ILF systems, investigating the different aspects including adversary and attack models to obtain meaningful data about the user’s identity, location, and time.
4.1 Adversary Models
@@
In general, user privacy means that no passive adversary (including a curious or malicious server) can determine the location of an honest user after intercepting all protocol messages. Server privacy, however, protects the server’s data from being compromised by malicious users using location queries [43]. To discuss threat models on privacy preservation in indoor localization, we need to explain privacy assumptions on how entities can behave in IPS. Four assumptions are taken into consideration which are explained in the following subsection.
4.1.1 Fully trusted
@@
4.1.2 Semi-trusted (honest-but-curious)
@@ 4.1.3 Fully malicious
4.1.4 Uunilateral-malicious
Introduced in [43], the unilateral-malicious setting is weaker than the fully malicious setting but stronger than the conventional Semi-trusted setting. We formulate the malicious behaviors specific to user sessions in the unilateral-malicious setting, such as manipulating Wi-Fi fingerprints and disclosing locations. We expect the server to act fairly honestly. In other words, while the server may be interested in learning the location of a user, it should run the protocol instance honestly to offer a high-quality service. Since a server providing inaccurate or unreliable location services would likely be noticed by users and lead them to stop using the service, the risk of deliberate poor service is lower. In other words, if the server consistently fails to show correct locations, users would abandon it, reducing the incentive for such attacks. However, the server faces more difficulty in detecting malicious behavior by users.
Understanding the difference between various settings lies in the level of trust and cooperation expected from the participants, as well as the defenses and security measures needed to ensure privacy and integrity in the face of potential adversarial behaviors and threats.
Figure 6: Adversary and attack models overview in ILF systems.
4.2 Attack Models
@@
4.2.1 Location Privacy and Data Privacy Attacks
Location privacy revolves around the potential threats to an individual’s privacy resulting from the collection, use, and disclosure of their location information. Targeted advertising, monitoring daily habits, surveillance, stalking and harassment, discriminatory uses, and unauthorized access are examples of location privacy concerns. Overall, location privacy is a subcategory of data privacy that is disparate in importance. By this means, from the perspective of the user, different locations have different privacy requirements. For instance, the majority of people worry significantly more about keeping their houses and workplaces private than they do about revealing where they visited in a shopping center. Despite differing levels of data sensitivity, the dangers associated with location privacy remain a substantial concern [39].
The major attack model on IPS is introduced in [44], which is followed by most papers. As shown in Fig. 7, this attack model categorizes the threats into the Location Privacy Attacks and Data Privacy Attacks. In both of them, location and data can be attacked in two ways as follows [44]:
@@ 4.2.1 Location Privacy and Data Privacy Attacks
Figure 7: Location privacy and data privacy attacks in ILF systems.
@@
Note that another categorization with the same principle is applied by using terms user-side privacy attack and server-side privacy attack [45]. The localization server and the APs are regarded as adversaries in breaking the user’s location privacy. The user, on the other hand, is seen as an adversary to compromise the server’s database security. The user must not reveal her location to untrusted parties, and the service provider should also keep its WiFi fingerprint database from unauthorized leaking. For example, a malicious user may download the database and sell it for profit.
4.2.2 Protocol-based attack
@@
In this attack, the malicious actor uses cryptographic methods to extract data from encrypted data or discovers misconfiguration in data transmission protocols between user and server. Specifically in the chosen fingerprint attack [45] or known location attack [43], the user chooses special fingerprints, such as all-zeros or single-one fingerprints, to compromise the whole server’s database.
4.2.3 Inference attack
An inference attack involves an adversary deducing sensitive information about individuals or systems by analyzing seemingly harmless or non-sensitive data. These attacks exploit patterns and relationships within the data to draw conclusions that may not be explicitly available. In the context of location-based services, inference attacks can reveal a person’s activities, habits, or even identity by analyzing patterns in their location data over time. Recent work in [46] demonstrates how inference attacks in indoor location models can exploit wireless signal data, such as Received Signal Strength Indicator (RSSI) values, to infer a device’s location without requiring physical access to the environment. By analyzing temporal signal strength patterns and employing techniques like Reinforcement Learning (RL), attackers can model device movements and deduce location information, posing significant privacy risks to indoor localization systems. One specific type of inference attack is a membership inference attack (MIA), where an adversary attempts to determine whether a particular individual’s data was used in training a machine learning model [47]. This form of attack has also been studied in indoor location models [48], where attackers analyze subtle cues from model outputs to infer whether specific location data was part of the training set, potentially compromising user privacy.
4.2.4 Semantic attack
@@
In a semantic attack, the objective of the attacker is to obtain specific semantic details about an individual. Location data inherently contains semantic information, and the disclosure of such information can unveil personal details and behavioral tendencies of users. Within indoor location datasets, an adversary can deduce the purpose of a user’s visits to particular locations by monitoring their movements or extended stays [49]. For example, in a hospital, the adversary may discern whether the user is visiting a specific doctor, and in a shopping center, they can identify specific shops where the user lingers. Additionally, this information might reveal whether the user is searching for or purchasing particular products in a shopping center, consulting a specific specialist in a hospital, or meeting a particular professor at a university.
4.2.5 Probability-based attack